OpenStudy (kevin8767):

Solve for a. 7a - 2b = 5a + b

OpenStudy (anonymous):

7a -5a = b+2b 2a = 3b a = 3/2*b = (3b)/2

OpenStudy (kevin8767):

i dont get this problem with the - sign in the middle

OpenStudy (kevin8767):

= sign